OpenStudy (trojanpoem):

\[1) \lim_{x \rightarrow 0} \frac{ \sin(\tan(5x)) }{ 4x }\] \[2) \lim_{x \rightarrow 0} \frac{ \tan(\sin(7x))}{ 3x }\] \[3) \lim_{x \rightarrow 0} \sin(7x(3+\frac{ 2 }{ x }))\] • Write the step, Thanks.

OpenStudy (anonymous):

\(1. \lim\limits_{x\to0}: \cfrac{\sin(\tan(5x))}{4x} \left[\frac00\right] = \cfrac{\cos(\tan(5x))\cdot\sec^2(5x)\cdot5}{4} = \cfrac{1\cdot1\cdot5}{4} = \cfrac54\) The result could alternatively be _motivated_ as folows: Near x=0, both sin(x) and tan(x) are roughly x, so sin(tan(5x)) is roughly 5x, and 5x/4x = 5/4.

OpenStudy (anonymous):

The \(\left[\frac00\right]\) means that we will use l'Hôpital's rule.

OpenStudy (trojanpoem):

Yeah, but sin(tan(5x)) -> cos(tan(5x))⋅sec2(5x)⋅5 What did you do to get that ?

OpenStudy (anonymous):

I used the chain rule.

OpenStudy (trojanpoem):

That's it thanks :D.

OpenStudy (anonymous):

No problem!

OpenStudy (trojanpoem):

Do you know the answer of the last one ?

OpenStudy (trojanpoem):

The second is almost the same as the first.

OpenStudy (anonymous):

Does \(\sin7x\,(3+\frac2x)\) mean \((\sin7x)(3+\frac2x)\) or \(\sin(7x\,(3+\frac2x))\)?

OpenStudy (trojanpoem):

The second one

OpenStudy (trojanpoem):

sin(7x(3+2x))

OpenStudy (anonymous):

\[\lim\limits_{x\to0}: \sin(7x\,(3+\frac2x)) = \sin(21x+14) = \sin(14)\]

OpenStudy (trojanpoem):

I found that it equals to sin(14) but isn't it strange i really never got something like that in limits

OpenStudy (anonymous):

It happens sometimes. A month ago wolframalpha evaluated this for me: \(\lim\limits_{x\to\infty} \left(\cfrac{(x+i)^2}{x^2+1}\right)^x = e^{2i}\).

OpenStudy (trojanpoem):

A nice problem , indeed.

OpenStudy (anonymous):

Wow, I just realized how to evaluate it! I'll show you :)

OpenStudy (trojanpoem):

Lol , Is it too long like that ? 0_0

OpenStudy (anonymous):

I'm using one definition of \(e^x\) saying that \(e^x = \lim\limits_{n\to\infty} (1+x/n)^n\). \[\begin{split} \lim\limits_{x\to\infty} \left(\cfrac{(x+i)^2}{x^2+1}\right)^x &= \lim\limits_{x\to\infty} \left(\cfrac{(x+i)^2/{x^2}}{(x^2+1)/{x^2}}\right)^x\\ &= \lim\limits_{x\to\infty} \left(\cfrac{(1+i/x)^2}{1+1/{x^2}}\right)^x\\ &= \lim\limits_{x\to\infty} \frac{(1+i/x)^{2x}}{(1+1/{x^2})^x}\\ &= \frac{\lim\limits_{x\to\infty}(1+i/x)^{2x}}{\lim\limits_{x\to\infty}(1+1/{x^2})^x}\\ \end{split}\] The numerator is \(\qquad\lim\limits_{x\to\infty}((1+i/x)^x)^2 = (e^x)^2 = e^{2x}\) and the denominator is \(\qquad\lim\limits_{x\to\infty}{(1+(1/x)/x})^x = \lim\limits_{x\to\infty} e^{1/x} = 1\) (This part could be made more formal.) so the quotient is \(e^{2x}\).

OpenStudy (trojanpoem):

limx→∞e1/x=1 (This part could be made more formal.) so the quotient is e2x. I can't understand this part the rest is fine to me.

OpenStudy (anonymous):

Sorry the \((e^x)^2\) etc should have an \(i\) instead of \(x\).

OpenStudy (anonymous):

I'll try to do it formally.

OpenStudy (trojanpoem):

"Sorry the (ex)2 etc should have an i instead of x." I didn't notice this mistake too because of the x in the saying. What's the hardest limit problem ? I wanna test myself.

OpenStudy (anonymous):

Well you could try to evaluate that denominator because I can't seem to figure it out.

OpenStudy (trojanpoem):

OII I will try.

OpenStudy (anonymous):

Any clues mathmate?

OpenStudy (trojanpoem):

Still working on it.

OpenStudy (anonymous):

I'll check back later.

OpenStudy (trojanpoem):

Okie.

OpenStudy (anonymous):

Here is how my teacher did it: \[\qquad\begin{split} L &= \lim\limits_{x\to\infty}{(1+1/x^2})^x \\ &= \lim\limits_{x\to\infty}{e^{\ln\left((1+1/x^2)^x\right)}} \\ &= \lim\limits_{x\to\infty}{e^{x\ln(1+1/x^2)}} \\ &= e^y\; (\text{where } y = \lim\limits_{x\to\infty}{x\ln(1+1/x^2)}) \\ &\qquad\begin{split} y &= \lim\limits_{x\to\infty} \frac{\ln(1+1/x^2)}{1/x} {\small\left[\frac00\right]}\\ &= \lim\limits_{x\to\infty} \frac{\left(\frac{-2/x^3}{1+1/x^2}\right)}{-1/x^2}\\ &= \lim\limits_{x\to\infty} \frac{2}{x+1/x}\\ &= 0 \end{split}\\ L &= e^0 = 1 \end{split}\]